# Confidential Serverless Computing

### v1

@@ 8 Evaluation
We evaluate Hacher by analyzing its end-to-end performance (§ 8.2),the performance of its operations (§ 8.3), its resource efficiency (§ 8.4), its communication network performance (§ 8.5), and its scale-out capabilities (§ 8.6).
@@ 8.1 Experimental Setup
Variants. We use the baselines summarized below..
@@
| Hacher | Our Hacher system |
@@ 8.2 End-to-end Performance
Methodology. We run the SeBS server in each variant and co-locate the client on the same host. Hacher does not include the time of measurement calculation for a fair comparison with the other baselines. We run each experiment five times and report the average.
End-to-end latency. Figure 9 presents the end-to-end client latency of each SeBS function for each variant. In the cold start case, Hacher (Lukewarm) achieves 93.44% speedup over CVM (SEV-SNP) on average, and is 83.99%, 51.04%, 32.85% faster than VM (KVM-Linux), Containers (Kata), and LibOS (Gramine). Even in the “cold”est case (i.e., no loaded zygote), Hacher is on average 85.10% faster than the CVM (SEV-SNP) baseline and 63.62% faster than the VM. When compared to the Containers (Kata), and LibOS (Gramine), Hacher incurs 11.25% and 52.57% performance overhead, respectively. In the warm start case, the performance difference among variants becomes smaller as the existing function environment is reused. Still, Hacher is 14.97%, 7.53% and 1.04% faster than the CVM, VM and Containers (Kata) case respectively while it shows a 15.63% performance overhead relative the LibOS (Gramine). Hacher’s performance gains mainly stem from its efficient function startup, run-to-completion execution model that minimizes VMEXITs, and in-memory filesystem. Hacher exhibits higher latency if the input/output size is large due to data transfer overhead between different VMPL levels (e.g., image-recognition has 100MB input) (§ 8.3).
Invocation latency. Figure 10 shows the cumulative distribution function (CDF) of invocation latency under cold starts, which primarily affect the invocation latency. Hacher (Lukewarm) has p50/p99 (50th and 99th percentile) latencies of 0.019/1.06 s. It achieves ms-scale p50 startup latency, comparable to the latest (non-confidential) fork-based serverless systems [25, 52, 113, 51, 94]. On the other hand, Hacher (cold) has 1.57/12.34 s latency, which is at the levels of Containers (1.93/2.10 s) and much lower than the CVM (21.25/22.10 s) and the VM (7.5/8.0 s) variants. It also exhibits a larger standard deviation (3.51 s), which originates from the time needed to load the zygote and the variably-sized functions (§ 8.3).
@@ 8.3 Performance Analysis
Boot time analysis. Figure 1 presents the boot time breakdown. For VM-based variants, the majority of the boot time is consumed by VMM (QEMU) initialization and guest OS startup. The CVM incurs additional overhead due to extra management tasks (e.g., measurement of initial state, memory validation) [109]. Although Hacher’s environment setup takes 4.05 s longer than the CVM (without memory preallocation (§ 8.3.1)) because of the added setup time for the monitor, Hacher’s cold start only takes 2.38 s by eliminating the CVM initialization time. Hacher’s lukewarm start further optimizes the boot time by employing a fork-based start (10.3 ms).
Runtime processing cost. Further, we analyze the runtime processing of Hacher, focusing on zygote and trustlet creation, and input/output transfer. Each process involves transferring data between the guest OS and the trusted monitor, followed by updating the VMPL level and the trustlet’s page tables. As Figure 10 illustrates, the zygote creation is responsible for most of the runtime initialization. The current prototype bundles required libraries in the base image (e.g., numpy, pytorch), increasing the zygote size, ranging from 60 MB to 691 MB (image-recognition) in SeBS. However, zygote creation occurs only during cold starts, and the CoW mechanism enables efficient memory sharing among trustlets using the same zygote image (§ 8.4).
@@
Measurement cost. Figure 2 shows the breakdown of the time taken to calculate the SHA512 measurement of an empty Python function. The calculation time is proportional to the data size. In the cold start case, the measurements include those of the trusted monitor, as well as the zygote (60 MB), which takes around 1.1 s. The current prototype does not support CPU acceleration for SHA, justifying the lower performance. However, for the lukewarm start, Hacher only needs to measure the function, input, and output (typically less than 4KB), dropping the calculation time below ms.
@@ 8.4 Resource Efficiency
Memory footprint. We evaluate the memory footprint of Hacher while running SeBS functions. Figure 10 presents the result. “CoW shared” represents the memory shared with CoW among trustlets and their base zygote, whereas “Non-shared” denotes the memory exclusively allocated for a trustlet during execution. Although the amount of allocated memory (“Non-shared”) varies depending on the workload, on average, trustlets share 170 MB of memory with the base zygote, reducing the total system memory consumption.
Function density. Memory sharing enables Hacher to pack more functions per node. We analyze the memory consumption for increasing numbers of concurrently-running empty Python functions, shown in Figure 2. We enable kernel same-page merging (KSM) [44] on the host to de-duplicate identical memory pages. For Hacher, the total memory is calculated as the base zygote shared memory (147 MB) plus 60 KB of non-CoW-shared memory per function. As KSM cannot deduplicate encrypted memory, CVM memory consumption is proportional to the number of functions, leading to a higher per-function memory overhead. Further, the number of CVMs in a single node is limited to the number of encryption keys (509 in our environment). In contrast, Hacher achieves significantly lower memory usage, realizing high function density.
@@ 8.5 Communication Analysis
Methodology. To evaluate Hacher’s inter-function communication, we measure the communication time of invoking multiple functions in a chain by varying the data size and the number of functions in the chain. All functions are deployed on the same host. In Hacher, each function creates its communication endpoint for input and output, while other baselines use TCP networking.
Communication latency. First, we show the communication latency between two functions under various data sizes. Figure 1 shows the result. This result indicates that Hacher achieves lower latency than CVMs (1.4-27), as well as traditional VMs and Containers (Kata), across message sizes, thanks to the zero-copy-based data objects.
Latencies of function chains. Next, we show the latencies of function chains. As shown in Figure 11, Hacher outperforms all baselines across all chain lengths, achieving 16.7-30.2, 11.8-20.9, and 15.2-30.1 lower communication latency than CVMs, Containers, and traditional VMs, respectively. Hacher’s performance advantage grows with chain length, demonstrating the efficiency of its data-centric I/O for inter-function communication compared to TCP-based networking used by other approaches, highlighting Hacher’s suitability for complex workloads with function chaining.
@@ 8.6 Scale-out Performance
Methodology. We develop a simulator modeling real-world serverless deployments with a central scheduler and multiple nodes, each having fixed execution slots and LRU caches for warm functions. The scheduler processes trace data chronologically, prioritizing nodes already caching the requested function. If none are available, it assigns requests to free nodes or queues them. Function execution duration gets increased based on its boot type (cold/lukewarm/warm), using timings sampled from our microbenchmarks (Figure 1). The lukewarm boot, unique to Hacher, occurs when a node caches another function from the same application, simulating zygote sharing. We use Azure Functions production traces [160, 139] sampled with InVitro [149], comprising 4k functions with 4.1 million invocations over 30 minutes.
Invocation latency and per-function slowdown. We run the simulator with 100 nodes, each having 32 execution slots and a cache size of 32. Figure 12 shows the CDF of invocation latency: at p50/p99, CVM and VM experience delays of 489/881 s and 1.3/33 s, respectively, while Hacher achieves just 5 ms/1.5 s. Figure 12 reports the per-function slowdown, which increases as function execution time decreases and invocation latency dominates. At p50/p99, CVM slowdowns reach 2,682/391,257, VM slowdowns are 13.98/8,658, while, in contrast, Hacher’s slowdowns are merely 1.02/7.06.
Node scalability. We also evaluate the invocation latency for various numbers of nodes. Figure 12 shows that, compared to the other baselines, Hacher achieves lower latency with a smaller number of nodes. For example, Hacher has 5 ms p99 latency with 100 nodes, while, even with 150 nodes, the CVM’s p99 latency is 50 s, and the VM’s is 12 s. These results highlight Hacher’s effectiveness in a large-scale environment.
@@ 9 Related Work
Serverless computing. Serverless computing has transformed cloud deployment. Open-source frameworks (OpenWhisk [47], OpenLambda [61], OpenFaaS [46]) and major cloud providers (AWS Lambda [138], Azure Functions [106], Google Cloud Functions [58]) offer widely used architectures. Recent research [78, 129, 79, 52, 160, 125, 141, 5, 113, 140, 25] mainly focuses on improving serverless isolation, performance and efficiency, while Hacher targets the security challenges and leverages TEEs to form a lightweight confidential computing system for secure serverless deployments.
@@
Confidential computing. Confidential computing, based on hardware TEEs [16, 9, 75, 17, 18, 74, 88], is widely adopted for protecting data and code in untrusted clouds. TEE technologies can be broadly categorized into process-based (e.g., Intel SGX [74]) and VM-based TEEs (e.g., AMD SEV-SNP [9], Intel TDX [75], ARM CCA [17]). Extensive recent research [115, 59, 162, 4, 11, 86, 55, 154, 126, 63, 56, 110, 131, 95, 150, 155, 90, 152, 158] aims to minimize TCB, and provide intra-VM isolation, among others.
@@
Specifically, Veil [4], and NestedSGX [152] deploy secure services in CVMs using SEV-SNP VMPL. Erebor [158] adopts intra-kernel privilege isolation to provide sandboxing in CVMs. Unlike these, Hacher targets serverless deployments, and leverages CVM partitioning for minimal TCB, fast boot, and efficient function communication via shared memory.
@@
Confidential serverless computing. Many research works aim for secure serverless computing architectures using confidential computing [120, 7, 92, 137, 82, 24, 54, 147, 116]. ServerlessCoCo [137] comprehensively analyzes overheads in CVM-based serverless deployments. Plugin Enclaves [92] proposes a hardware-based approach with Intel SGX for efficient confidential serverless computing. Cryonics [82] reduces startup times using snapshot-based SGX enclaves. In contrast, Hacher materializes confidential serverless computing by providing a CVM-based lightweight system with easily attestable and deployable trusted processes.
@@
Lightweight virtualization. Lightweight virtualization techniques [3, 41, 85, 26, 38, 86, 83, 100, 98] aim to reduce the VM overheads while maintaining isolation. Firecracker [3] introduces microVMs for serverless workloads with minimal boot times. Kata Containers [41] deploy lightweight VMs with container-level isolation in standard VMs. Unikraft [85] creates specialized unikernels for applications. Gramine LibOS [26, 38, 86] runs unmodified apps in SGX enclaves and TDX VMs. Hacher leverages these concepts with a minimal LibOS for its trusted processes, enabling low boot and attestation times, and seamless application deployment.
@@
Networking for confidential computing. TEE networking is typically slow due to the added overhead of encryption and data copying with multiple context switches [90, 109, 153]. Several works optimize their processing overhead with polling at the cost of increased CPU usage [145, 146, 20]. Bifrost [90] optimizes CVM network processing by removing redundant bounce buffer copying while preserving protocol safety guarantees. Compared to them, Hacher achieves efficient inter-function communication through its data-centric I/O architecture while ensuring isolation.
@@ 10 Conclusion
In this paper, we present Hacher, a lightweight confidential serverless computing system. Hacher leverages intra-CVM partitioning mechanisms and efficiently consolidates security-critical functionalities into a compact, privileged trusted monitor, resulting in a small TCB. Through its zygote mechanism, Hacher achieves fast boot times while optimizing memory management and reducing the communication cost between serverless functions via its data-centric I/O architecture. Further, it provides a formally verified differential attestation mechanism to ensure end-to-end trust with reduced attestation times. Overall, Hacher introduces low-performance overheads while exposing a minimal attack surface, making it ideal for security- and latency-sensitive serverless deployments.
Acknowledgements
This work was supported in part by an ERC Starting Grant (ID: 101077577) and the Chips Joint Undertaking (JU), European Union (EU) HORIZON-JU-IA, under grant agreement No. 101140087 (SMARTY), the Intel Trustworthy Data Center of the Future (TDCoF), and Google Research Grants. The authors acknowledge the financial support by the Federal Ministry of Education and Research of Germany in the programme of "Souverän. Digital. Vernetzt.". Joint project 6G-life, project identification number: 16KISK002.

### v3

@@ 8 Evaluation
We evaluate Wallet by analyzing its end-to-end performance (§ 8.2),the performance of its operations (§ 8.3), its resource efficiency (§ 8.4), its communication network performance (§ 8.5), and its scale-out capabilities (§ 8.6).
@@ 8.1 Experimental Setup
Variants. We use the baselines summarized below.
@@
| Wallet | Our Wallet system |
Workloads. For performance evaluation, we use Python functions from the SeBS benchmark suite [34]. The current Wallet prototype lacks external network and fork() support; therefore, we exclude “uploader”, “crud-api” (requires external networking), and “video-processing” (requires running ffmpeg as a standalone process). Additionally, we conduct a simulation-based study with Azure Functions production traces [165, 142].
@@ 8.2 End-to-end Performance
Methodology. We run the SeBS server in each variant and co-locate the client on the same host. Wallet does not include the time of measurement calculation for a fair comparison with the other baselines. We run each experiment five times and report the average.
End-to-end latency. Figure 9 presents the end-to-end client latency of each SeBS function for each variant. In the cold start case, Wallet (Lukewarm) achieves 93.44% speedup over CVM (SEV-SNP) on average, and is 83.99%, 51.04%, 32.85% faster than VM (KVM-Linux), Containers (Kata), and LibOS (Gramine). Even in the “cold”est case (i.e., no loaded zygote), Wallet is on average 85.10% faster than the CVM (SEV-SNP) baseline and 63.62% faster than the VM. When compared to the Containers (Kata), and LibOS (Gramine), Wallet incurs 11.25% and 52.57% performance overhead, respectively. In the warm start case, the performance difference among variants becomes smaller as the existing function environment is reused. Still, Wallet is 14.97%, 7.53% and 1.04% faster than the CVM, VM and Containers (Kata) case respectively while it shows a 15.63% performance overhead relative the LibOS (Gramine). Wallet’s performance gains mainly stem from its efficient function startup, run-to-completion execution model that minimizes VMEXITs, and in-memory filesystem. Wallet exhibits higher latency if the input/output size is large due to data transfer overhead between different VMPL levels (e.g., image-recognition has 100MB input) (§ 8.3).
Invocation latency. Figure 10 shows the cumulative distribution function (CDF) of invocation latency under cold starts, which primarily affect the invocation latency. Wallet (Lukewarm) has p50/p99 (50th and 99th percentile) latencies of 0.019/1.06 s. It achieves ms-scale p50 startup latency, comparable to the latest (non-confidential) fork-based serverless systems [26, 53, 116, 52, 96]. On the other hand, Wallet (cold) has 1.57/12.34 s latency, which is at the levels of Containers (1.93/2.10 s) and much lower than the CVM (21.25/22.10 s) and the VM (7.5/8.0 s) variants. It also exhibits a larger standard deviation (3.51 s), which originates from the time needed to load the zygote and the variably-sized functions (§ 8.3).
@@ 8.3 Performance Analysis
Boot time analysis. Figure 1 presents the boot time breakdown. For VM-based variants, the majority of the boot time is consumed by VMM (QEMU) initialization and guest OS startup. The CVM incurs additional overhead due to extra management tasks (e.g., measurement of initial state, memory validation) [112]. Although Wallet’s environment setup takes 4.05 s longer than the CVM (without memory preallocation (§ 8.3.1)) because of the added setup time for the monitor, Wallet’s cold start only takes 2.38 s by eliminating the CVM initialization time. Wallet’s lukewarm start further optimizes the boot time by employing a fork-based start (10.3 ms).
Runtime processing cost. Further, we analyze the runtime processing of Wallet, focusing on zygote and trustlet creation, and input/output transfer. Each process involves transferring data between the guest OS and the trusted monitor, followed by updating the VMPL level and the trustlet’s page tables. As Figure 10 illustrates, the zygote creation is responsible for most of the runtime initialization. The current prototype bundles required libraries in the base image (e.g., numpy, pytorch), increasing the zygote size, ranging from 60 MB to 691 MB (image-recognition) in SeBS. However, zygote creation occurs only during cold starts, and the CoW mechanism enables efficient memory sharing among trustlets using the same zygote image (§ 8.4).
@@
Measurement cost. Figure 2 shows the breakdown of the time taken to calculate the SHA512 measurement of an empty Python function. The calculation time is proportional to the data size. In the cold start case, the measurements include those of the trusted monitor, as well as the zygote (60 MB), which takes around 1.1 s. The current prototype does not support CPU acceleration for SHA, justifying the lower performance. However, for the lukewarm start, Wallet only needs to measure the function, input, and output (typically less than 4KB), dropping the calculation time below ms.
@@ 8.4 Resource Efficiency
Memory footprint. We evaluate the memory footprint of Wallet while running SeBS functions. Figure 10 presents the result. “CoW shared” represents the memory shared with CoW among trustlets and their base zygote, whereas “Non-shared” denotes the memory exclusively allocated for a trustlet during execution. Although the amount of allocated memory (“Non-shared”) varies depending on the workload, on average, trustlets share 170 MB of memory with the base zygote, reducing the total system memory consumption.
Function density. Memory sharing enables Wallet to pack more functions per node. We analyze the memory consumption for increasing numbers of concurrently-running empty Python functions, shown in Figure 2. We enable kernel same-page merging (KSM) [45] on the host to de-duplicate identical memory pages. For Wallet, the total memory is calculated as the base zygote shared memory (147 MB) plus 60 KB of non-CoW-shared memory per function. As KSM cannot deduplicate encrypted memory, CVM memory consumption is proportional to the number of functions, leading to a higher per-function memory overhead. Further, the number of CVMs in a single node is limited to the number of encryption keys (509 in our environment). In contrast, Wallet achieves significantly lower memory usage, realizing high function density.
@@ 8.5 Communication Analysis
Methodology. To evaluate Wallet’s inter-function communication, we measure the communication time of invoking multiple functions in a chain by varying the data size and the number of functions in the chain. All functions are deployed on the same host. In Wallet, each function creates its communication endpoint for input and output, while other baselines use TCP networking.
Communication latency. First, we show the communication latency between two functions under various data sizes. Figure 1 shows the result. This result indicates that Wallet achieves lower latency than CVMs (1.4-27), as well as traditional VMs and Containers (Kata), across message sizes, thanks to the zero-copy-based data objects.
Latencies of function chains. Next, we show the latencies of function chains. As shown in Figure 11, Wallet outperforms all baselines across all chain lengths, achieving 16.7-30.2, 11.8-20.9, and 15.2-30.1 lower communication latency than CVMs, Containers, and traditional VMs, respectively. Wallet’s performance advantage grows with chain length, demonstrating the efficiency of its data-centric I/O for inter-function communication compared to TCP-based networking used by other approaches, highlighting Wallet’s suitability for complex workloads with function chaining.
@@ 8.6 Scale-out Performance
Methodology. We develop a simulator modeling real-world serverless deployments with a central scheduler and multiple nodes, each having fixed execution slots and LRU caches for warm functions. The scheduler processes trace data chronologically, prioritizing nodes already caching the requested function. If none are available, it assigns requests to free nodes or queues them. Function execution duration gets increased based on its boot type (cold/lukewarm/warm), using timings sampled from our microbenchmarks (Figure 1). The lukewarm boot, unique to Wallet, occurs when a node caches another function from the same application, simulating zygote sharing. We use Azure Functions production traces [165, 142] sampled with InVitro [154], comprising 4k functions with 4.1 million invocations over 30 minutes.
Invocation latency and per-function slowdown. We run the simulator with 100 nodes, each having 32 execution slots and a cache size of 32. Figure 12 shows the CDF of invocation latency: at p50/p99, CVM and VM experience delays of 489/881 s and 1.3/33 s, respectively, while Wallet achieves just 5 ms/1.5 s. Figure 12 reports the per-function slowdown, which increases as function execution time decreases and invocation latency dominates. At p50/p99, CVM slowdowns reach 2,682/391,257, VM slowdowns are 13.98/8,658, while, in contrast, Wallet’s slowdowns are merely 1.02/7.06.
Node scalability. We also evaluate the invocation latency for various numbers of nodes. Figure 12 shows that, compared to the other baselines, Wallet achieves lower latency with a smaller number of nodes. For example, Wallet has 5 ms p99 latency with 100 nodes, while, even with 150 nodes, the CVM’s p99 latency is 50 s, and the VM’s is 12 s. These results highlight Wallet’s effectiveness in a large-scale environment.
@@ 9 Related Work
Serverless computing. Serverless computing has transformed cloud deployment. Open-source frameworks (OpenWhisk [48], OpenLambda [63], OpenFaaS [47]) and major cloud providers (AWS Lambda [141], Azure Functions [108], Google Cloud Functions [60]) offer widely used architectures. Recent research [80, 132, 81, 53, 165, 128, 146, 5, 116, 144, 26] mainly focuses on improving serverless isolation, performance and efficiency, while Wallet targets the security challenges and leverages TEEs to form a lightweight confidential computing system for secure serverless deployments.
@@
Confidential computing. Confidential computing, based on hardware TEEs [16, 9, 77, 17, 18, 76, 90], is widely adopted for protecting data and code in untrusted clouds. TEE technologies can be broadly categorized into process-based (e.g., Intel SGX [76]) and VM-based TEEs (e.g., AMD SEV-SNP [9], Intel TDX [77], ARM CCA [17]). Extensive recent research [118, 61, 167, 4, 11, 88, 56, 159, 129, 65, 58, 113, 134, 97, 155, 160, 92, 157, 163, 145, 57, 21] aims to minimize TCB, and provide intra-VM isolation, secure data at rest, among others.
@@
Specifically, Veil [4], and NestedSGX [157] deploy secure services in CVMs using SEV-SNP VMPL. Erebor [163] adopts intra-kernel privilege isolation to provide sandboxing in CVMs. Unlike these, Wallet targets serverless deployments, and leverages CVM partitioning for minimal TCB, fast boot, and efficient function communication via shared memory.
@@
Confidential serverless computing. Many research works aim for secure serverless computing architectures using confidential computing [123, 7, 94, 140, 84, 25, 55, 152, 119]. ServerlessCoCo [140] comprehensively analyzes overheads in CVM-based serverless deployments. Plugin Enclaves [94] proposes a hardware-based approach with Intel SGX for efficient confidential serverless computing. Cryonics [84] reduces startup times using snapshot-based SGX enclaves. In contrast, Wallet materializes confidential serverless computing by providing a CVM-based lightweight system with easily attestable and deployable trusted processes.
Concurrently with our work, COFUNC [143] proposes a library OS-based confidential containers tailored for serverless functions. COFUNC employs a microkernel architecture, running multiple containers with the same CVM while enforcing isolation. Contrary to Wallet, COFUNC offloads container management tasks to the serverless runtime that runs on the host. On the other hand, Wallet allows for the reuse of the existing OS and serverless framework within the CVM, leveraging its nested confidential execution architecture.
@@
Lightweight virtualization. Lightweight virtualization techniques [3, 42, 87, 27, 39, 88, 85, 102, 100] aim to reduce the VM overheads while maintaining isolation. Firecracker [3] introduces microVMs for serverless workloads. Kata Containers [42] deploy lightweight VMs for container applications, whereas Unikraft [87] creates specialized unikernels for applications. Gramine LibOS [27, 39, 88] runs unmodified apps in SGX enclaves and TDX VMs [88]. On the other hand, Wallet leverages these concepts with a minimal Gramine LibOS instance for each of its trusted processes, enabling low boot and attestation times, and seamless application deployment. Unlike Gramine-TDX, Wallet’s LibOS is only running in a certain privilege level within the CVM and does not have system-wide privileges. Further, in contrast to Gramine’s single-tenant model with its existing backends, Wallet provides a framework for managing multiple instances within the same CVM, enabling multitenancy and optimized inter-function communication via data objects among trusted processes while maintaining function instance protection.
@@
Networking/IO for confidential computing. TEE networking is typically slow due to the added overhead of encryption and data copying with multiple context switches [92, 112, 158] Several works optimize their processing overhead with polling at the cost of increased CPU usage [150, 151, 20, 21]. Compared to them, Wallet achieves efficient inter-function communication through its data-centric shared-memory-based I/O architecture while ensuring isolation.
@@ 10 Conclusion
In this paper, we present Wallet, a lightweight confidential serverless computing system. Wallet leverages intra-CVM partitioning mechanisms and efficiently consolidates security-critical functionalities into a compact, privileged trusted monitor, resulting in a small TCB. Through its zygote mechanism, Wallet achieves fast boot times while optimizing memory management and reducing the communication cost between serverless functions via its data-centric I/O architecture. Further, it provides a formally verified differential attestation mechanism to ensure end-to-end trust with reduced attestation times. Overall, Wallet introduces low-performance overheads while exposing a minimal attack surface, making it ideal for security- and latency-sensitive serverless deployments.
Artifact. Wallet is publicly available at https://github.com/TUM-DSE/Wallet-VMPL along with its entire experimental setup.
Acknowledgments
We thank our shepherd, Prof. Anurag Khandelwal, and the anonymous reviewers for their helpful comments. We also thank Prof. Nuno Santos for valuable discussion and feedback on this work.
This work was supported in part by an ERC Starting Grant (ID: 101077577) and the Chips Joint Undertaking (JU), European Union (EU) HORIZON-JU-IA, under grant agreement No. 101140087 (SMARTY), the Intel Trustworthy Data Center of the Future (TDCoF), and Google Research Grants. The authors acknowledge the financial support by the Federal Ministry of Research, Technology and Space of Germany in the programme of “Souverän. Digital. Vernetzt.”. Joint project 6G-life, project identification number: 16KISK002.